OpenStudy (anonymous):

A triangle has sides of the square root of 2 and 3. Which could not be the length of the third side if it is a right triangle? A. The square root of 7 B. The square root of 11 C. The square root of 13 I do not know how to do these type of problems, so I would really like it if someone were to show me how to do them. :)

OpenStudy (anonymous):

whoch one do you think is corrcet

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@rosy_love

OpenStudy (anonymous):

why do you think b is you answer?

OpenStudy (anonymous):

idk how to do these problems so i just guessed

OpenStudy (deana99):

@raiders88 I just saw this

OpenStudy (anonymous):

:) i think you're corrcey

OpenStudy (welshfella):



OpenStudy (welshfella):

by pythagoras x^2 = (sqrt2)^2 + 3^2 = 2 + 9 = 11

OpenStudy (welshfella):

B

OpenStudy (welshfella):

in any right angled triangle the square of the hypotenuse (the longest side) = the sum of the squares of the other 2 sides

OpenStudy (welshfella):

so if you know the values of any 2 sides of a right angled triangle you use this (Pythagoras) theorem to find the measure of the third side.

OpenStudy (welshfella):

Note this only applies to a right angled triangle - not any triangle